Circular No. 8251

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003kj–2003kv P. Challis, Center for Astrophysics, on behalf of the ESSENCE project

(cf. IAUC 8237, 8021, 8034), reports the discovery of 13 supernovae on multiple R- and I-band images taken at the Blanco 4-m telescope during Nov. 15–30 by project members. Spectroscopy — obtained on Nov. 15–30 with the Very Large Telescope (+ FORS1), Gemini telescope (+ GMOS), and Magellan (+ IMACS and + LDSS2) reflectors — shows all except SN 2003kj (which is of type II) to be of type Ia. Included below are measured redshifts; each new supernova is within 1′′–2′′ of its host-galaxy center.

SN  2003 UT  α_{2000}  δ_{2000}  Mag. z
2003kj Nov. 15  1^{h}14^{m}36.^{s}56 − 0^{o}23′58′′.2 22.3 0.1
2003kk Nov. 15 23 25 36.06 − 9 31 44.7 20.3 0.2
2003kl Nov. 17  1 09 48.80 + 1 00 05.6 21.0 0.3
2003km Nov. 19  2 30 09.00 − 9 04 35.9 21.8 0.3
2003kn Nov. 21  2 09 15.55 − 3 35 41.4 21.3 0.2
2003ko Nov. 21  2 11 06.48 − 3 47 56.1 21.7 0.4
2003kp Nov. 21  2 31 02.64 − 8 39 50.8 22.1 0.6
2003kq Nov. 21  2 31 04.09 − 8 10 56.6 22.6 0.6
2003kr Nov. 21  2 31 20.96 − 8 36 14.2 22.0 0.4
2003ks Nov. 21  2 31 34.54 − 8 36 46.4 22.2 0.5
2003kt Nov. 21  2 33 47.01 − 8 36 22.1 22.8 0.6
2003ku Nov. 23  1 08 36.25 − 0 33 20.8 22.9 0.8
2003kv Nov. 27  2 09 42.52 − 3 46 48.6 23.4 0.8
2003 UN_{284}

R. L. Millis, Lowell Observatory, on behalf of the Deep Ecliptic Survey team, reports that the recently reported transneptunian object 2003 UN_{284} (cf. MPEC 2003-X20) is a binary system. Scrutiny by K. B. Clancy (Massachusetts Institute of Technology) of two images taken on Oct. 24 UT at the Kitt Peak 4-m Mayall Telescope (+ MOSAIC camera) revealed the two objects having a separation of 2′′.01 ± 0′′.11 and a broad-VR-bandpass-magnitude difference of 0.59 ± 0.21 mag, the fainter component lying in p.a. 305^{o} ± 3^{o} relative to the brighter component. Subsequent observations on Nov. 20 (two images) and Nov. 24 (two images) with the same instrumentation yielded corresponding values of 1′′.93 ± 0′′.10, 0.70 ± 0.15 mag, and p.a. 299^{o} ± 3^{o}. The brighter component has R ≈ 23.2. Additional astrometric and photometric observations of 2003 UN_{284} are encouraged.

2003 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green